OpenStudy (anonymous):

log base 2 of (x) = sqrt(x) solve for all x values where this is true.

OpenStudy (jdoe0001):

\(\huge log_{2}(x)=\sqrt{x}\implies 2^{\frac{x}{2}}=x\)

OpenStudy (anonymous):

I actually got close to this point. What I didn't figure out was how to solve for x once you get to the equation: \[2^{\sqrt{x}} = x\] (which is equivalent to what you solved for).

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

for one yours is correct, mine isn't

OpenStudy (jdoe0001):

\(\huge log_{2}(x)=\sqrt{x}\implies 2^{x^\frac{1}{2}}=x\implies 2^{\sqrt{x}}=x\) would be correct

OpenStudy (jdoe0001):

I can see.. offhand. that 4 works

OpenStudy (anonymous):

Really? I thought \[2^{x /2}\] was equivalent to \[2^{x^{1/2}}\] which is the same as \[2^{x / 2}\] since when a power is raised to a power you just multiply. Right?

OpenStudy (jdoe0001):

16 works too

OpenStudy (anonymous):

And yeah, i know that 4 and 16 work, but i going for how to solve these in a general sense. I tried to pick as simple of an example as I could come up with.

OpenStudy (jdoe0001):

heheh, well \(\bf \large x^{\frac{1}{2}}\ne \frac{x}{2}\)

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

oh, derp. yeah.

OpenStudy (jdoe0001):

well... hmmm there seem to be just a few values that will fit the bill for the base 2 to match up "x"

OpenStudy (anonymous):

there are only two. I know that much

OpenStudy (jdoe0001):

well... I agree =)